Beer is fermenting in a steel tank that is 5m high. This tank has an access hatch on the side 15 cm above the bottom, for maintenance purposes. The hatch is elliptical with a width of 65cm and a hight of 35 cm. Two bolts of 10mm diameter are used to hold the hatch in place. Determine the stress in each bolt when the level of the beer is 4m above the top of the hatch. Assume the beer to have a density of 650kg/m^3 and the bolts to be horizontally on either side of the centre of pressure. At what depth is the centre of pressure?


1
Expert's answer
2021-06-04T16:29:03-0400

"P=\rho hg=650*5*9.81=31850"

Area of elliptical shape = "\pi r_1r_2=\pi *\frac{65}{100}* \frac{55}{100}=0.71471"

Force "= PA =31850*0.71471=22763.58766N"

Area of each bolt "=\pi r^2=\pi*(\frac{10}{200})^2=0.00007"

Force on the bolt "=P*A=31850*0.00007=2.2295N"

H="h_1+h_2=5m+\frac{15*2}{100}+4m=9.3m"